40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 2 11:37:38 2019, gautam, Update, ALS, EY IR ALS Assay IMG_7780.JPGEy_MM_20190802.pdf
    Reply  Tue Aug 6 23:09:20 2019, gautam, Update, ALS, EY table work 
       Reply  Fri Aug 9 16:37:39 2019, gautam, Update, ALS, More EY table work IMG_7791.JPGfiberCleaning.pdf
Message ID: 14823     Entry time: Fri Aug 2 11:37:38 2019     Reply to this: 14835
Author: gautam 
Type: Update 
Category: ALS 
Subject: EY IR ALS Assay 

Summary:

I'd like to confirm that the IR ALS scheme will work for locking. The X-arm performance so far has been encouraging. I want to repeat the characterization for the Y arm. So I inspected the layout on the EY table, and made a list of characterization tasks. The current EY beam routing is difficult to work with, and it will definitely benefit from a re-do. However, I don't know how much time I want to spend re-doing it, so for a start, I will just try and couple some amount of light into a fiber and bring it to the PSL table, and see what noise performance I get.

Details:

Attachment #1: Photo of the current beam layout. The powers indicated were measured with the Ophir power meter.

  • I measure an SHG conversion efficiency of 0.87 %/W, which is considerably lower than the ~3.7%/W that is theoretically expected, and 1.5%/W that is realized at EX.
  • Of the 0.5 mW of green light that is generated, I measure ~0.375 mW at the viewport into the EY chamber. So there is ~25 % loss in the green beam path on the EY table. Seems high to me.
  • The previous solution of coupling IR light into the fiber realized at EY was to use the SHG leakage IR beam. While there isn't a measurement showing that this dirty beam is noisier than a cleaner pickoff, I'd like to adopt the solution used at EX, which is to use the leakage beam from the first steering mirror in the NPRO beam path. This will allow better mode-matching and polarization control of the beam being coupled into the fiber, which at least in principle, translates to less phase noise.
  • However - the beam layout at the EY table offers much less freedom to work with this idea than EX. A constraint is the clamp that secures the enclosure to the optical table, labelled in the photo. Further behind it, the green steering optics occupy all available space. A more comprehensive photo of the EY table can be found here.
  • Off the top of my head, I don't see any other good open spots on the EY table where we could couple IR light into the fiber. 
  • One other change I'd like to make is to replace the first steering mirror after the NPRO head, which is currently a Y1 HR mirror, with a R=99% BS. This will make it easier to control the amount of power coupled into the fiber, which is something we'd like.

Attachment #2: A candidate mode-matching solution, given the constraints outlined above. It isn't great, with only 85% modematching even theoretically possible. The lenses required are also pretty fast lenses. But I think it's the best possible without a complete overhaul of the EY layout. I'm still waiting for the lens kit to arrive, but as soon as they get here, I will start this work.

Characterization tasks:

  • Characterize SHG at EY [done 7/28]
  • Characterize gPDH at EY (loop TFs, improve MM, PDH discriminant, check the polarization)
  • Couple IR light into fiber with good MM at EY [done with 36% MM 8/9]
  • Clean fiber at EY, and at the PSL table [done 8/9]
  • Make the PSL + Y IR beat [done 8/9]
  • Noise budget
Attachment 1: IMG_7780.JPG  1.339 MB  Uploaded Fri Aug 2 12:59:24 2019  | Hide | Hide all
IMG_7780.JPG
Attachment 2: Ey_MM_20190802.pdf  59 kB  Uploaded Fri Aug 2 13:06:21 2019  | Hide | Hide all
Ey_MM_20190802.pdf
ELOG V3.1.3-